There are 5 questions on a multiple choice exam, each with five possible answers. If a student guesses on all five questions, what is the probability the student gets three questions correct? Express your answer as a percent rounded to the nearest tenth.
A) 16.9%
B) 5.1%
C) 20.5%
D) 8.7%

Answer by stanbon(75887)

There are 5 questions on a multiple choice exam, each with five possible answers. If a student guesses on all five questions, what is the probability the student gets three questions correct? Express your answer as a percent rounded to the nearest tenth.
Binomial Problem with n = 5 and p = 1/5
-----
P(x=3) = 5C3*(1/5)^3*(4/5)^2 = binompdf(5,1/5,3) = 5.1%
----------------
Cheers,
Stan H.
A) 16.9%
B) 5.1%
C) 20.5%
D) 8.7%